DO NOT ANSWER IF YOU ARE UNSKILLED IN THIS AREA!

Using 20 observations, the following regression output is obtained from estimating y = β0 + β1x + β2d + β3xd + ε.

Coefficients Standard Error t Stat p-value
Intercept 14.14 2.80 5.05 0.000
x 4.93 0.50 9.86 0.000
d −5.72 5.50 −1.04 0.314
xd 1.04 0.80 1.30 0.212


a. Compute yˆy^ for x = 10 and d = 1; then compute yˆy^ for x = 10 and d = 0. (Round intermediate calculations to at least 4 decimal places and final answers to 2 decimal places.)

x = 10 and d = 1
x = 10 and d = 0
Answer #1

a) x = 10, d = 1

So,

y = 14.14 + 4.93*10 - 5.72*1 + 1.04*10*1 = 68.12

b) x = 10, d = 0

So,

y = 14.14 + 4.93*10 = 63.44